OpenStudy (anonymous):

solve by addition 5x+3y=-1 x+y=1

OpenStudy (anonymous):

y = 3, x = -2

OpenStudy (anonymous):

If you want more than a mere answer without any explanation, here goes We have to eliminate a variable. To do so we need to multiply a constant Let's eliminate y Multiply second equation by -3, we get -3x-3y=-3 add that to first equation 5x+3y=-1 -3x-3y=-3 ----------- 2x=-4 x=-2 Find y by plugging in value of x